Consider cobalt-60 (⁶⁰CO) with a half-line of 5.25 years. How many years will it take for its activity to decrease to one-fourth of its original amount?

1
Expert's answer
2022-05-30T15:18:17-0400
"A=A_0*2^{-t\/t_{1\/2}}"

"\frac{1}{4}A_0=A_0*2^{-t/t_{1/2}}"

"t=2t_{1/2}=2*5.25\:\rm yr=10.5\:yr"